# An Encoding–Searching Separation Perspective on Bi-Encoder Neural Search

### v1

@@ Abstract
This paper reviews, analyzes, and proposes a new perspective on the bi-encoder architecture for neural search. While the bi-encoder architecture is widely used due to its simplicity and scalability at test time, it has some notable issues such as low performance on seen datasets and weak zero-shot performance on new datasets. In this paper, we analyze these issues and summarize two main critiques: the encoding information bottleneck problem and limitations of the basic assumption of embedding search. We then construct a thought experiment to logically analyze the encoding and searching operations and challenge the basic assumption of embedding search. Building on these observations, we propose a new perspective on the bi-encoder architecture called the encoding–searching separation perspective, which conceptually and practically separates the encoding and searching operations. This new perspective is applied to explain the root cause of the identified issues and discuss ways to mitigate the problems. Finally, we discuss the implications of the ideas underlying the new perspective, the design surface that it exposes and the potential research directions arising from it.
Keywords:
information retrieval neural search bi-encoder encoding searching perspectives.
@@ 2.3 Bi-Encoder Architecture
[…]
@@ 5.3 New Research Directions
  • Design Customized Models for Transfer Learning: This research direction aims to explore new configurations and design choices enabled by the encoding–searching separation perspective that are suitable for zero-shot search and transfer learning between search tasks, an important topic in the field of neural search.
@@ 6 Related Work
There have been several surveys on neural search such as [12] providing an introduction to neural search, [8] describing the emerging trend of using pretrained transformers for neural search. However, these surveys do not specifically review and analyze the identified issues of the bi-encoder architecture in this paper. The issue of low zero-shot search result has been discussed extensively in [18], which inspired us to look deeper into the problems and its root cause. The proposed perspective provides an explanation to this issue and suggests possible solutions.
@@
There are several extensions of the bi-encoder architecture such as ColBERT [7] that uses all tokens’ embeddings as encoding output and a specialized similarity function, Poly-encoder [4] that generates multiple embeddings for the query and the item, then uses a complex similarity function to search, which is costly at test time, because the search items need to be accessed one-by-one. These works have some similarity to the proposed perspective in that they have some functions on top of the encoding output. However, they do not explicitly view the model in the perspective of encoding–searching separation. In fact, the proposed perspective may be used to explain some properties of these models.
The proposed perspective shares some similarities with some old neural search methods predating transformers that work on fixed word embeddings [12]. However, they did not have access to context-rich generic encoding operation such as pretrained large language models, as a result, their perspectives did not consider the generic embedding as an advantage, and usually focused on developing complex similarity functions. Instead, we extensively analyze the generic encoding–specific searching setting to gain better understandings of modern bi-encoder neural search and explore potential improving directions.
@@ 7 Conclusion
In this paper, we analyzed and summarized the main critiques of the bi-encoder architecture, including the encoding information bottleneck problem and limitations of the basic assumption of embedding search. We conducted a thought experiment to challenge the basic assumption of embedding search and proposed the new encoding–searching separation perspective, which conceptually and practically separate the encoding and the searching operations.
Based on the new perspective, we tried to explain the root causes of the identified issues of the bi-encoder architecture and explored potential mitigations. Finally, we discussed the implications of the new perspective, the design surface it exposes, and potential research directions arising from it. Overall, our perspective contributes to the understanding of the bi-encoder architecture and suggest potential directions for advancing neural search research.
@@ 0.A.1 On Equivalence and Difference
It should be noted that the encoding–searching separation perspective still describes the original practical bi-encoder architecture in Eq. 2. In fact, we can show that they are mathematically equivalent as follows.
@@
$enc1^{\prime\prime}(\cdot)=$ $f_{search}\circ enc1^{\prime}(\cdot)$
=\ = $f_{search}\big{(}enc1^{\prime}(q)\big{)},$ (4)
@@
It should be noted that certain extensions of the bi-encoder architecture, which use more complex similarity functions, such as the poly-encoder model [4] and ColBERT [7], may look similar to the proposed perspective but are distinct from it. Firstly, practically, these models use multiple embeddings, and their complex similarity functions may require the item embeddings to be accessed one-by-one at test time, as is the case with poly-encoder. Secondly, and more importantly, conceptually, they do not separate the searching from the encoding operation and are thus still limited by the encoding-for-search assumption.

### v2

# Revisiting Bi-Encoder Neural Search: An Encoding–Searching Separation Perspective

@@ Abstract
This paper reviews, analyzes, and proposes a new perspective on the bi-encoder architecture for neural search. While the bi-encoder architecture is widely used due to its simplicity and scalability at test time, it has some notable issues such as low performance on seen datasets and weak zero-shot performance on new datasets. In this paper, we analyze these issues and summarize two main critiques: the encoding information bottleneck problem and limitations of the basic assumption of embedding search. We then construct a thought experiment to logically analyze the encoding and searching operations and challenge the basic assumptions of embedding search. Building on these observations, we propose a new perspective on the bi-encoder architecture called the encoding–searching separation perspective, which conceptually and practically separates the encoding and searching operations. This framework is applied to explain the root cause of existing issues and suggest mitigation strategies, potentially lowering training costs and improving retrieval performance. Finally, we discuss the broader implications of the ideas underlying this perspective, the new design surface it exposes, and potential research directions arising from it.
@@ 2.3 Bi-Encoder Architecture
where $\langle\cdot,\cdot\rangle$ denotes the inner product between the embedding vectors. Note that the inner product is fast not just because it is simple, but because the item embeddings can be indexed in advance, instead of requiring computation one-by-one as in other approaches.
@@ 5.3 New Research Directions
  • Design Customized Models for Transfer Learning: This research direction aims to explore new configurations enabled by the encoding–searching separation perspective that are suitable for zero-shot search and transfer learning between search tasks, an important topic in the field of neural search.
@@ 6 Related Work
There have been several surveys on neural search such as [12] providing an introduction to neural search, [8] describing the emerging trend of using pretrained transformers for neural search. However, these surveys do not specifically review and analyze the identified issues of the bi-encoder architecture in this paper. The issue of low zero-shot search result has been discussed extensively in [18], which inspired us to look deeper into the problems and its root cause. The proposed perspective provides an explanation to this issue and suggests potential solutions.
@@
There are several extensions of the bi-encoder architecture such as ColBERT [7] that uses all tokens’ embeddings as encoding output and a specialized similarity function, Poly-encoder [4] that generates multiple embeddings for the query and the item, then uses a complex similarity function to search, which is costly at test time, because the search items need to be accessed one-by-one. These works have some similarity to the proposed perspective in that they have some functions on top of the encoding output. However, they do not explicitly view the model in the perspective of encoding–searching separation. In fact, our approach may be used to explain some properties of these models.
Our approach shares some similarities with some old neural search methods predating transformers that work on fixed word embeddings [12]. However, they did not have access to context-rich generic encoding operation such as pretrained large language models, as a result, their perspectives did not consider the generic embedding as an advantage, and usually focused on developing complex similarity functions. Instead, we extensively analyze the generic encoding–specific searching setting to gain better understandings of modern bi-encoder neural search and explore potential improving directions.
@@ 7 Conclusion
In this paper, we analyzed and summarized the main critiques of the bi-encoder architecture, including the encoding information bottleneck problem and limitations of the basic assumption of embedding search. Through a thought experiment, we challenged the basic assumption of embedding search and proposed the new encoding–searching separation perspective, which conceptually and practically separate the encoding and the searching operations. Based on the new perspective, we explained the root causes of the identified issues of the bi-encoder architecture and explored potential mitigations. Finally, we discussed the broader implications of the new perspective, the design surface it exposes, and potential research directions arising from it.
While this paper introduces a new perspective on bi-encoder architectures and provides theoretical insights, it currently lacks extensive empirical experiments across diverse datasets. Demonstrating practical advantages, such as improved retrieval performance and reduced training costs, requires rigorous empirical evaluation. To address this, we plan to conduct experiments using MS MARCO [1] for supervised training and BEIR [18] for zero-shot evaluation, comparing our approach with standard bi-encoders in terms of accuracy, computational efficiency, and the trade-off between accuracy and compute. A particularly promising direction is training an ensemble of lightweight searching modules on top of precomputed frozen embeddings, enabling fast and efficient training. Investigating how far these modules can bridge the encoding gap and how their composition contributes to generalization across different zero-shot datasets would provide valuable insights. We hope this work lays the foundation for future research, inviting further exploration into the interplay between encoding and searching to advance neural retrieval models.
Appendix
@@ 0.A.1 On Equivalence and Difference
Note that the encoding–searching separation perspective still describes the original practical bi-encoder architecture in Eq. 2. In fact, we can show that they are mathematically equivalent as follows.
@@
$enc1^{\prime\prime}(\cdot)=\$$f_{search}\circ enc1^{\prime}(\cdot)$
[…]
@@
Certain extensions of the bi-encoder architecture, which use more complex similarity functions, such as the poly-encoder model [4] and ColBERT [7], may look similar to the proposed perspective but are distinct from it. Firstly, practically, these models use multiple embeddings, and their complex similarity functions may require the item embeddings to be accessed one-by-one at test time, as is the case with poly-encoder. Secondly, and more importantly, conceptually, they do not separate the searching from the encoding operation and are thus still limited by the encoding-for-search assumption.